# Algorithmic and High-Frequency Trading Problems for Semi-Markov and Hawkes Jump-Diffusion Models

### v1

@@ 3 Stochastic Optimal Control Trading Problems: Acquisition and Liquidation
SOC theory is utilized in a variety of different continuous-time problems, in order to optimize the performance of particular model. Some of the first historical examples in finance were first developed in the 1970s, with the first seminal paper by Merton, (1969) being devoted to the portfolio allocation problem. The results of this model were then extended by many others, such as Zariphopoulou, (1989) to optimal investment-consumption models, Davis and Norman, (1990) to portfolio selection with transaction costs, and Oksendal and Sulem, (2002) to the portfolio selection problem with fixed and variable transaction costs. In this paper, we will focus on applying the SOC framework to algorithmic and HFT problems, where one can find a whole range of examples in Cartea et al., (2015), one of which we will solve with our new price processes from section 2.
In general, the agent in our examples would either like to acquire/buy or liquidate/sell a large number of units (shares, lots, etc) over a certain time horizon, say from $[0,T]$, where it would be sub-optimal to acquire/liquidate the targeted number of units instantly. These types of scenarios occur regularly and particularly when the total amount to be acquired or liquidated is much larger than the available liquidity in the LOB. The general setup of these types of trading problems involves using stochastic optimal control theory, where the goal is to maximise/minimize an objective function by acting in an optimal manner. Thus, the agent can control a certain action it takes and this is referred to as the control. The control variable in the case of either the acquisition or liquidation problem we study is the speed/rate of trading. The optimal control is then driven by the dynamics of the asset price and many other processes (deterministic or stochastic) deemed relevant to the problem. The key stochastic processes in these types of trading problems, as described in Cartea et al., (2015), are as follows,
  • $\nu=(\nu_{t})_{\{0\leq t\leq T\}}$ is the trading speed/rate, the speed at which the agent is acquiring units. This is the variable the agent can control.
@@
  • $S^{\nu}=(S_{t}^{\nu})_{\{0\leq t\leq T\}}$ is the midprice process. This is also affected by the speed of the agent’s trades and so is the agent’s controlled stock price process.
@@
  • $C^{\nu}=(C_{t}^{\nu})_{\{0\leq t\leq T\}}$ is the agents cash process resulting from the execution strategy.
@@
    Here, $g,\eta_{SM},\eta_{HP},\varsigma_{SM},\varsigma_{HP},\sigma$, $\bar{\sigma}^{2}_{SM}$, $\bar{\sigma}^{2}_{HP}$ and $W$ are as defined in section 2. The permanent price impact function, $g$, is positive in an acquisition problem as the agents actions create a upward drift in the price and negative in a liquidation problem as then the agents actions create a downward drift in the price. Next, we will rewrite this midprice process one more time so that we have just one price process for both cases and we will proceed with this formulation when solving the acquisition and liquidation problems later on. This price process will be defined as,
@@ 3.1 Acquisition Problem
In this trading problem, as explained in Cartea et al., (2015), the agent’s objective is to acquire $\mathfrak{N}$ units over a trading horizon $T$, with a cap on the price at which she acquires units equal to $S_{max}$. The agent will stop trading if the agent has acquired $\mathfrak{N}$ units, the terminal time $T$ is reached or the midprice $S_{t}$ (given by (17) with a positive drift term) has reached the price cap $S_{max}$. To formalise this, define a stopping time as
@@
Here, remember the coefficients $\eta$, $\bar{\sigma}$ and $\varsigma$ are representing either the cases for $\eta_{SM}$, $\bar{\sigma}_{SM}$ and $\varsigma_{SM}$, or $\eta_{HP}$, $\bar{\sigma}_{HP}$ and $\varsigma_{HP}$, from section 2.1 and 2.2, respectively. One can see from the terminal and boundary conditions, that once the stopping time (see equation (20)) is reached, the agent will purchase the remainder of the units. Also, one can see that once the agent has purchased the $\mathfrak{N}$ units initially targeted, the agent stops trading and there is no penalty, hence, the boundary condition $H(t,S,0)=0$ where $y=0$.
@@
Next, as in Cartea et al., (2015) we reduce the dimensions in order to be able to solve this problem numerically. In order to do so, we first set $b=0$ using their same argument that the permanent impact of walking the LOB tends to be relatively small compared to the temporary impact. Then, we can use the following ansatz,
@@
This PDE now has a new terminal and boundary condition defined as,
@@
Here, we can see that the midprice path taken is paramount in computing the level of inventory the agent has acquired at any given time. In section 5, we will show some examples, via strategy simulations, how the optimal strategy for different midprice paths plays out.
@@ 3.2 Liquidation Problem
In this trading problem, the agent’s objective is to liquidate $\mathfrak{N}$ units again over a trading horizon $[0,T]$, where now there is a floor on the price at which the agent liquidates units equal to $S_{min}$. The setup here is similar to the acquisition problem in section 3.1, except from a liquidation standpoint. The agent will stop trading if the agent has liquidated $\mathfrak{N}$ units, the terminal time $T$ is reached or the midprice $S_{t}$ has hit the price floor $S_{min}$. To formalise this, we define a stopping time as
@@
Here, remember the coefficients $\eta$, $\bar{\sigma}$ and $\varsigma$ are representing either the cases for $\eta_{SM}$, $\bar{\sigma}_{SM}$ and $\varsigma_{SM}$, or $\eta_{HP}$, $\bar{\sigma}_{HP}$ and $\varsigma_{HP}$, from section 2.1 and 2.2, respectively. One can see from the terminal and boundary conditions, that once the stopping time is reached, the agent will liquidate the remainder of the units. Also, one can see that once the agent has liquidated the $\mathfrak{N}$ units initially targeted, the agent stops trading and there is no penalty, hence the boundary condition $H(t,S,0)=0$ where $q=0$.
Then, as in section 3.1, one can obtain the optimal trading speed, by using the mathematical procedure called completing the squares on the inside of the minimization part of equation (38). And so, the optimal trading speed in feedback form can be defined as,
@@
Next, as in Cartea et al., (2015) and section 3.1, we reduce the dimensions in order to be able to solve this problem numerically. In order to do so, we first set $b=0$ using the same argument as in section 3.1 that the permanent impact of walking the LOB tends to be relatively small compared to the temporary impact. Then, we can use the following ansatz,
@@
This PDE now has a new terminal and boundary condition defined as,
@@
Lastly, as in Cartea et al., (2015) for the acquisition problem, we would like to portray how the current status of the inventory process Q can be obtained in terms of the path of the midprice process by computing $Q_{t}$, the process which determines how much inventory is left to liquidate. Here, $Q_{t}$ satisfies the SDE,
@@ 4 Numerical Solution: Acquisition and Liquidation
In this section, we will discuss the numerical solution to the acquisition and liquidations problems from section 3.1 and 3.2, respectively. We solve the PDEs in equation (29) and (43) numerically using an Implicit-Explicit(IMEX) finite difference scheme, similarly to how they solve an acquisition problem in Cartea et al., (2015) for the pure diffusion case. The solutions are placed on a $[0,T]\times[S_{min},S_{max}]$ grid, which is the domain. Then, for $n=0,1,...,N\in\mathbb{N}$, define $\Delta t=T/N$ such that $t_{n}=n\Delta t$ and for $i=0,1,...,M\in\mathbb{N}$, define $\Delta S=(S_{max}-S_{min})/M$ such that $S_{i}=S_{min}+i\Delta S$. The grids are divided into equally spaced nodes of distance $\Delta t$ and $\Delta S$, with mesh points $(n\Delta t,S_{min}+i\Delta S)$. We are concerned with the values of $h(t,S)$ i.e., the speed of trading per unit of inventory left to acquire/liquidate. The numerical scheme solves the $\partial_{t}h$ and $\partial_{ss}h$ partial derivatives in equations (29) and (43) implicitly using standard methods and explicitly for the quadratic term as in Cartea et al., (2015).
@@ 4.1 Acquisition Problem Solution
Note that one must also account for the boundary conditions, at $i=0$ and $i=M$, and the terminal condition, at $n=N$. These are as defined in equation (30)-(31) and the condition above where a lower boundary was set at $S_{min}$.
In Figure 1, we show a set of subplots which display contour plots portraying how the solution changes as a function of time and the financial asset’s price, over different values of $\bar{\sigma}$ and $\varsigma$. Time here is from $t=0$ to the maturity time $t=T=1$, and the midprice (the space variable) is from $S_{0}=20$ to $S_{max}=20.20$. We specifically focus on this part of the solution as this is where it varies the most. Right of each subplot shows the color scheme for the values of the solution that are displayed. These solution values vary from 00 to $0.01$, where $\alpha=0.01$ represents the value of the terminal condition and the boundary condition at $T$ and $S_{max}$, respectively, as defined in equations (28)-(29). We can see that the agents speed of acquiring units increases as we get closer to the terminal time. This is because the agent is trying to avoid the terminal penalty cost that they must pay to purchase any remaining units at maturity. Furthermore, the speed of trading also increases as the asset’s price approaches the limit price, $S_{max}$, again to avoid the other terminal penalty. This trading process was formalized in the stopping time in equation (20).
@@
In the top left subplot of Figure 1, we show the solution for the case where both $\bar{\sigma}=0$ and $\varsigma=0$, which would represent the case for benchmark models in Cartea et al., (2015). In other words, this subplot would portray the same solution were our price process to follow the pure diffusion dynamics given in equation (1). In the top right subplot, we show the solution for the case where $\bar{\sigma}=0.1$ and $\varsigma=0$, which would portray the same solution were the pricing model to follow the price process from Cartea et al., (2015) combined with the price processes developed in Roldan Contreras and Swishchuk, (2022) and Roldan Contreras, (2023). Recall that the latter models approximate the price process, via a diffusion approximation, for the cases where the dynamics approximate the semi-Markov or Hawkes processes. The bottom two subplots, where we increased the value of the parameter $\varsigma$, then show the solutions for the cases where the price process follows our more general jump-diffusion model given in equation (17).
@@
In each consecutive subplot, from left to right, we vary either $\bar{\sigma}$ or $\varsigma$ by $0.1$ to show the effect that the Semi-Markov and Hawkes process parts can have on the optimal solution. One can see from the subplots that as we slowly increment either of these parameters, the speed of trading increases at earlier time steps at lower asset prices. This makes sense, intuitively, as the higher either of these coefficients becomes, the more overall volatility the solution predicts. Thus, in these scenarios, it would then become more optimal to acquire units of the asset quicker based on our model. In this particular type of trading problem, where the agent is acquiring units, it’s mainly price jumps that go up which can significantly change the optimal solution as the agent will want to acquire units of the asset a lot quicker then. In the case where the asset price would jump down, this would be highly favourable to the agent, and their speed of acquiring any remaining units of the asset would then mostly depend on time.
@@
To see the remaining input parameters for this model, please see Table 1 below. Most of these parameters were kept similar to the simulation results in Cartea et al., (2015) for comparison purposes, where they provide some details for their choices. We did, however, make some changes. Firstly, we increased the spatial size of the grid to better illustrate how our new parameters can influence a problem of this nature. To do this, we increased $S_{max}$ from $20.1$ to $20.2$ and we decreased $S_{min}$ from $19.5$ to $18.2$. Secondly, we decreased the value of $\phi$ from $0.001$ to $0.00001$, the running inventory penalty, so that our solution mostly focuses on the effect of the Semi-Markov and Hawkes Process parts of the model. This parameter must be nonzero for our strategy simulations later, which is why we decided to make it very small instead of setting it to zero. Next, as explained in Cartea et al., (2015), the formula $\alpha/\kappa$ represents the maximum trading speed, which is why $\alpha=0.01$. The choice of volatility is based on the formula $\sigma=0.1(T)^{-0.5}$, where $T=1$ in our case, which represents an annualized realized volatility of $0.1/20\times\sqrt{255}\approx 8\%$, where $255$ is the average number of trading days in a given year. These are also mostly the same parameters we used for our strategy simulations and that Cartea et al., (2015) used in their strategy simulations, which we will show in section 5.

### v2

@@ 3 Stochastic Optimal Control Trading Problems: Acquisition and Liquidation
SOC theory is utilized in a variety of different continuous-time problems to optimize the performance of a particular model. Some of the first historical examples in finance were developed in the 1970s in the seminal paper by Merton, (1969) which is devoted to the portfolio allocation problem. The results of this model were then extended by many others, such as Zariphopoulou, (1989) to optimal investment-consumption models, Davis and Norman, (1990) to portfolio selection with transaction costs, and Oksendal and Sulem, (2002) to the portfolio selection problem with fixed and variable transaction costs. In this paper, we will focus on applying the SOC framework to algorithmic and HFT problems, where a whole range of examples can be found in Cartea et al., (2015), one of which we will solve with our new price processes from Section 2.
In general, the agent in our examples would either like to acquire/buy or liquidate/sell a large number of units (shares, lots, etc) over a certain time horizon, say from $[0,T]$, where it would be sub-optimal to acquire/liquidate the targeted number of units instantly. These types of trading scenarios occur regularly and particularly when the total amount to be acquired or liquidated is much larger than the available liquidity in the LOB. The general setup of these types of trading problems involves using stochastic optimal control theory, where the goal is to maximize/minimize an objective function by acting in an optimal manner. Thus, the agent can regulate the specific set of actions it can take, which is referred to as the control. The control variable in the case of either the acquisition or liquidation problem we study is the speed/rate of trading. The optimal control is then driven by the dynamics of the asset price and many other processes (deterministic or stochastic) deemed relevant to the problem. The key stochastic processes in these types of trading problems, as described in Cartea et al., (2015), are as follows,
  • $\nu=(\nu_{t})_{\{0\leq t\leq T\}}$ is the trading speed/rate, the speed at which the agent is buying or selling units of an asset. This is the variable that the agent can control.
@@
  • $S^{\nu}=(S_{t}^{\nu})_{\{0\leq t\leq T\}}$ is the midprice process. This is also affected by the speed of the agent’s trades and so is the agent’s controlled price process.
@@
  • $C^{\nu}=(C_{t}^{\nu})_{\{0\leq t\leq T\}}$ is the agent’s cash process resulting from the execution strategy.
@@
    Here, $g,\eta_{SM},\eta_{HP},\varsigma_{SM},\varsigma_{HP},\sigma$, $\bar{\sigma}^{2}_{SM}$, $\bar{\sigma}^{2}_{HP}$ and $W$ are as defined in Section 2. The permanent price impact function, $g$, is positive in an acquisition problem as the actions of the agent create an upward drift in the price and negative in a liquidation problem as the actions of the agents create a downward drift in the price. Some of the foundational work on price impact models includes Bertsimas and Lo, (1998), which focuses on optimizing trade execution costs given a price impact function. In addition, Almgren and Chriss, (2001) and Colaneri et al., (2020) introduce permanent and temporary market impact as forms of transaction costs in their trading models. The latter also addresses an optimal liquidation problem, which is one of the trading problems examined in this study. Next, we will rewrite this midprice process one more time so that we have just one price process for both cases, and we will proceed with this formulation when solving the acquisition and liquidation problems later on. This price process will be defined as,
@@ 3.1 Acquisition Problem
In this trading problem, as explained in Cartea et al., (2015), the agent’s objective is to acquire $\mathfrak{N}$ units over a trading horizon $T$, with a cap on the price at which the agent acquires units equal to $S_{max}$. The agent will stop trading if the agent has acquired $\mathfrak{N}$ units, the terminal time $T$ is reached, or the midprice $S_{t}$ (given by Equation (17)) with a positive drift term) has reached the price cap $S_{max}$. To formalize this, define a stopping time as
@@
Here, remember the coefficients $\eta$, $\bar{\sigma}$ and $\varsigma$ are representing either the cases for $\eta_{SM}$, $\bar{\sigma}_{SM}$ and $\varsigma_{SM}$, or $\eta_{HP}$, $\bar{\sigma}_{HP}$ and $\varsigma_{HP}$, from Equations (3) and (9), respectively. One can see from the terminal and boundary conditions, that once the stopping time (see Equation (20)) is reached, the agent will purchase the remainder of the units. Also, one can see that once the agent has purchased the $\mathfrak{N}$ units initially targeted, the agent stops trading and there is no penalty, hence, the boundary condition $H(t,S,0)=0$ where $y=0$.
@@
Next, as in Cartea et al., (2015) we reduce the dimensions in order to be able to solve this problem numerically. In order to do so, we first set $b=0$ using their same argument that the permanent impact of walking the LOB tends to be relatively small compared to the temporary impact. This point is also emphasized in Almgren and Chriss, (2001). Although this simplifies the modeling approach, one slight drawback is that the model could underestimate any long-term market impacts that may occur. And so, we can use the following ansatz,
@@
The existence of solutions to this type of PDE follow from the same case as in Cartea et al., (2015) in the general case, which can be proven to be guaranteed under the so-called verification theorem following along the lines of many popular works such as Pham, (2009) and Yong and Zhou, (2012). This PDE now has a new terminal and boundary condition defined as,
@@
Here, we can see that the midprice path taken is paramount in computing the level of inventory the agent has acquired at any given time. In Section 5, we will show some examples, through strategy simulations, how the optimal strategy performs over different midprice paths.
@@ 3.2 Liquidation Problem
In this trading problem, the agent’s objective is to liquidate $\mathfrak{N}$ units again over a trading horizon $[0,T]$, where now there is a floor on the price at which the agent liquidates units equal to $S_{min}$. This setup is similar to the acquisition problem in Section 3.1 but viewed from a liquidation perspective. The agent will stop trading if the agent has liquidated $\mathfrak{N}$ units, the terminal time $T$ is reached, or the midprice $S_{t}$ has hit the price floor $S_{min}$. To formalise this, we define a stopping time as
@@
Here, remember the coefficients $\eta$, $\bar{\sigma}$ and $\varsigma$ are representing either the cases for $\eta_{SM}$, $\bar{\sigma}_{SM}$ and $\varsigma_{SM}$, or $\eta_{HP}$, $\bar{\sigma}_{HP}$ and $\varsigma_{HP}$, from Equations (3) and (9), respectively. One can see from the terminal and boundary conditions, that once the stopping time is reached, the agent will liquidate the remainder of the units. Also, one can see that once the agent has liquidated the $\mathfrak{N}$ units initially targeted, the agent stops trading and there is no penalty, hence the boundary condition $H(t,S,0)=0$ where $q=0$.
Then, as in Section 3.1, the optimal trading speed can be obtained by using the mathematical procedure called completing the squares on the inside of the minimization part of Equation (38). And so, the optimal trading speed in feedback form can be defined as,
@@
Following Cartea et al., (2015) and Section 3.1, we simplify the dimensionality to facilitate a numerical solution to this problem. To achieve this, we first set $b=0$ using the same argument as in Section 3.1 that the permanent impact of walking the LOB tends to be relatively small compared to the temporary impact. Then, we can use the following ansatz,
@@
The existence of solutions to this PDE follows along similar lines as mentioned in Section 3.1 for the acquisition problem, where the general case for the Liquidation problem is also studied in Cartea et al., (2015). This PDE now has a new terminal and boundary condition defined as,
@@
Lastly, as in Cartea et al., (2015) for the acquisition problem, we would like to portray how the current status of the inventory process, $Q$, can be obtained in terms of the path of the midprice process by computing $Q_{t}$, the process that determines how much inventory is left to liquidate. Here, $Q_{t}$ satisfies the SDE,
@@ 4 Numerical Solution: Acquisition and Liquidation
In this Section, we will discuss the numerical solution to the acquisition and liquidation problems from Section 3.1 and 3.2, respectively. Here, we adopt an asset-specific approach, focusing on Microsoft (MSFT), with LOB data sourced from LOBSTER, (2025). They offer free data on one day, June 21st 2012, on this asset. Building on this, we utilize the previously calibrated parameters from Swishchuk and Huffman, (2020) and Roldan Contreras, (2023), assuming that Equation (17) follows the Hawkes dynamics described in Equation (9). In addition, we conducted a sensitivity analysis to examine how varying these parameters affects the results. We solve the PDEs in Equation (29) and (43) numerically using an Implicit-Explicit(IMEX) finite difference scheme, similarly to how they solve an acquisition problem in Cartea et al., (2015) for the pure diffusion case. The solutions are placed on a $[0,T]\times[S_{min},S_{max}]$ grid, which is the domain. Then, for $n=0,1,...,N\in\mathbb{N}$, define $\Delta t=T/N$ such that $t_{n}=n\Delta t$ and for $i=0,1,...,M\in\mathbb{N}$, define $\Delta S=(S_{max}-S_{min})/M$ such that $S_{i}=S_{min}+i\Delta S$. The grids are divided into equally spaced nodes of distance $\Delta t$ and $\Delta S$, with mesh points $(n\Delta t,S_{min}+i\Delta S)$. We are concerned with the values of $h(t,S)$ i.e., the speed of trading per unit of inventory left to acquire/liquidate. The numerical scheme solves the $\partial_{t}h$ and $\partial_{ss}h$ partial derivatives in Equation (29) and (43) implicitly using standard methods and explicitly for the quadratic term as in Cartea et al., (2015).
@@ 4.1 Acquisition Problem Solution
Note that one must also account for the boundary conditions, at $i=0$ and $i=M$, and the terminal condition, at $n=N$. These were defined in Equation (30)-(31) and the condition above where a lower boundary was set at $S_{min}$.
In Figure 2, we show a set of subplots which display contour plots portraying how the solution changes as a function of time and the financial asset’s price, over different values of $\bar{\sigma}$ and $\varsigma$, where the calibrated parameter values are indicated in bold. Time here is from $t=0$ to the maturity time $t=T=1$, and the midprice (the space variable) is from $S_{0}=30.97$ to $S_{max}=31.1$. We specifically focus on this part of the solution as this is where it varies the most. Right of each subplot shows the color scheme for the values of the solution that are displayed. These solution values vary from 00 to $0.01$, where $\alpha=0.01$ represents the value of the terminal condition and the boundary condition at $T$ and $S_{max}$, respectively, as defined in Equations (30)-(31). We can see that the agents speed of acquiring units increases as we get closer to the terminal time. This is because the agent is trying to avoid the terminal penalty cost that they must pay to purchase any remaining units at maturity. Furthermore, the speed of trading also increases as the asset’s price approaches the limit price, $S_{max}$, again to avoid the other terminal penalty. This trading process was formalized in the stopping time in Equation (20).
@@
In the top left subplot of Figure 2, we show the solution for the case where both $\bar{\sigma}=0$ and $\varsigma=0$, which would represent the case for benchmark models in Cartea et al., (2015). In other words, this subplot would portray the same solution were our price process to follow the pure diffusion dynamics given in Equation (1). In the top right subplot, we show the solution for the case where $\bar{\sigma}=0.01598$ as calibrated and $\varsigma=0$, which shows the same solution if the pricing model follows the price process from Cartea et al., (2015) combined with the price processes developed in Roldan Contreras and Swishchuk, (2022) and Roldan Contreras, (2023). Recall that the latter models approximate the price process, via a diffusion approximation, for the cases where the dynamics approximate the semi-Markov or Hawkes processes. The two bottom subplots, where we increased the value of the parameter $\varsigma$ to its calibrated parameter value and then a higher value, show the solutions for the cases where the price process follows our more general jump-diffusion model given in Equation (17).
@@
In each consecutive subplot, from left to right, we vary either $\bar{\sigma}$ or $\varsigma$ to show the effect that the Semi-Markov or Hawkes process parts could have on the optimal solution. One can see from the subplots that as we slowly increment either of these parameters, the speed of trading increases at earlier time steps at lower asset prices. This makes sense intuitively, as the higher either of these coefficients becomes, the more overall volatility the solution predicts. Thus, in these scenarios, it would become more optimal to acquire units of the asset faster on the basis of our model. In this particular type of trading problem, where the agent is acquiring units, it is mainly price jumps that go up which can significantly change the optimal solution as the agent will want to acquire units of the asset a lot quicker then. In the case where the asset price would jump down, this would be highly favorable to the agent, and their speed of acquiring any remaining units of the asset would then mostly depend on time.
@@
To see the remaining input parameters for this model, please see Table 1 below. Most of these parameters were kept similar to the simulation results in Cartea et al., (2015) for comparison purposes, where they provide some details of their choices. However, we did make some changes. Firstly, we increased the spatial size of the grid to better illustrate how our new parameters can influence a problem of this nature. To do this, we slightly increase the distance between $S_{0}$ and $S_{max}$ and decrease $S_{min}$ by a greater amount as this part of the solution is less relevant as lower prices reduce the problem to a time-dependent execution strategy. Secondly, we decreased the value of $\phi$ from $0.001$ to $0.00001$, the running inventory penalty, so that our solution mostly focuses on the effect of the Semi-Markov and Hawkes process parts of the model. This parameter must be nonzero for our strategy simulations later which is why we decided to make it very small instead of setting it to zero. Next, as explained in Cartea et al., (2015), the formula $\alpha/\kappa$ represents the maximum trading speed, which is why $\alpha=0.01$. The choice of volatility, which was calibrated independently on the MSFT data using the maximum likelihood estimator approach, was set to $\sigma=0.1041$. These are also mostly the same parameter values used for our strategy simulations and similar to the ones that Cartea et al., (2015) used in their strategy simulations, which we will show in Section 5.